OpenStudy (jordans_openstudy):

Write the point-slope form of the equation of a line with slope 2/5 that passes through the point (-3,-6)

OpenStudy (mathmale):

What does the point-slope form look like? \[y-y _{1}=m(x-x _{1})\] You are given x, y and b. Plug these values into the equation to find the specific equation for this situation.

OpenStudy (jordans_openstudy):

so would "b" be "m'?

OpenStudy (welshfella):

Yes

OpenStudy (greatlife44):

What do you think it is?

OpenStudy (mathmale):

Sorry to contradict anyone, but no, b is not m. b in this context is the y-intercept; m is the slope, and there's no way in which these are the same.

OpenStudy (jordans_openstudy):

so would 2/5 go where M is and then -3 for y1, and -6 for x1?

OpenStudy (greatlife44):

yeah exactly

OpenStudy (greatlife44):

Now put it all together

OpenStudy (jordans_openstudy):

what about x and y?

OpenStudy (greatlife44):

what's your formula again tell us what you're plugging those numbers into

OpenStudy (mathmale):

"Write the point-slope form of the equation of a line with slope 2/5 that passes through the point (-3,-6)." Here the slope, m, is (2/5). x_1 is -3. y_1 is -6. You need to subst. these 3 numerical values into the point-slope formula I gave your earlier.

OpenStudy (mathmale):

Please review this entire conversation before proceding.

OpenStudy (jordans_openstudy):

y-(-6)=2/5(x-(-3)) ?

OpenStudy (mathmale):

If you would please enclose "2/5" inside parentheses, that'd give you the perfect response (answer).

OpenStudy (jordans_openstudy):

y=2/5x-24/5

OpenStudy (jordans_openstudy):

thats what i got from a online calculator i used

OpenStudy (mathmale):

Again, please enclose that 2/5 within parentheses, so there's no doubt to anyone that you mean 2x/5 and not 2/(5x).

OpenStudy (mathmale):

I like 2x/5 and (2/5)x, but NOT 2/5x.

OpenStudy (jordans_openstudy):

so y=(2/5)x-(24/5)

OpenStudy (jordans_openstudy):

that would be the correct answer then ?

OpenStudy (mathmale):

Yes! Thank you for the medal.

OpenStudy (jordans_openstudy):

Not a problem! i appreciate the help.